# LoopSR: Looping Sim-and-Real for Lifelong Policy Adaptation of Legged Robots

### v1

@@ IV-B Sim-to-Real Evaluation
Overview. For real-world evaluations, we utilize the Unitree A1 robot [62] to facilitate real-world deployment. The pretraining process is identical to the sim-to-sim experiments, and the testing environment consists of diverse real-world terrains, as depicted in Fig. 4.
Data collected in the real world are transmitted to a host computer with a GTX 4090 GPU to execute the continual training. Trajectories are collected with a length of seconds each. After each batch of trajectories is collected, the agent is trained for an episode of iterations in the simulation. Here, we only update the onboard policy after episodes to simplify the data collection process. We demonstrate that this simplicity approximates the real workflow for the empirical consistency in simulation parameters predicted from trajectories collected by different policies.
The real-world observation includes proprioception , obtained or estimated from the onboard sensors and IMU. The action specifies target joint positions converted to joint torques via a PD controller with stiffness and damping set to and , respectively. The policy operates at Hz.
We compare our method to Origin in Sec. IV-A to highlight the improvement. Each evaluation includes trials to reduce stochasticity while obviating damage to robot hardware. The performance metrics include the time required to traverse through a specified path and the success rate in demanding terrains.
@@
| LoopSR(Ours) | |||
| Origin |
Results. The results in Fig. 5 and Tab. II demonstrate that LoopSR achieves superior performance across all tasks compared with Origin, and ensures better safety during operation. The results indicate that LoopSR enhances locomotion policy capacity and fulfills the purpose of lifelong policy adaptation.
@@ IV-D Ablation Study
We conduct the ablation study to verify each component of LoopSR. To determine the importance of each loss, we test the accuracy of the variants in identifying different terrains and robot parameters. We also provide the policy without a soft update in training to validate its importance. The variants are designed as:
LSR-w/o-con eliminates the constrastive loss with only the multihead predictor to learn the latent space.
LSR-w/o-AE dismisses the reconstruction loss in the training process and depends on the constrastive loss to drive learning.
LSR-w/o-su excludes the soft update in continual training. Results in Tab. III are the percentages of rewards from LSR-w/o-su to the rewards from LSR.
From Tab. III, LSR with all losses outperforms all counterparts to classify the terrains accurately. The comparable failure of LSR-w/o-con signifies the importance of the contrastive loss [53] in building a robust and interpretable latent space. The comparison between LSR and LSR-w/o-AE proves that the autoencoder structure is beneficial in extracting enough information from the trajectories. LSR-w/o-su drastically underperforms on demanding terrains, demonstrating the necessity of soft update.
| Domain | LSR | LSR-w/o-con | LSR-w/o-AE | LSR-w/o-su |

### v2

@@ IV-B Sim-to-Real Evaluation
Overview. For real-world evaluations, we utilize the Unitree A1 robot [56] to facilitate real-world deployment. Real-world observation of A1 only includes proprioception without visual input. The action specifies target joint positions converted to joint torques via a PD controller. The testing environment consists of diverse real-world terrains, as depicted in Fig. 4.
The pretraining phase is identical to the sim-to-sim ones. For the policy adaptation phase, data collected in the real world are transmitted to a host computer with a GTX 4090 GPU to execute the continual training. To be more specific, we run 30 loops in total, and in each loop, we collect a batch of trajectories, each with length of timesteps (equal to seconds). After each batch of trajectories is collected, the agent is trained for iterations in the simulation. To simplify the data collection process, here we update the onboard policy every loops (equals simulation iterations), which we find has little impact on the adaptive performance.
Domain | LoopSR | Origin | RMA |
Stair Up | |||
Stair Down | |||
Grass | |||
Slope Down | |||
Slope Up |
We compare LoopSR to Origin (DreamWaQ) and RMA baselines to highlight the improvement. Each evaluation includes trials to reduce stochasticity while obviating damage to robot hardware. The performance metrics include the time required to traverse a specified path and the success rate in demanding terrains.
@@
LoopSR | |||
Origin | |||
RMA |
Results. The results in Tab. II and Tab. III demonstrate that LoopSR achieves superior performance across all tasks compared with zero-shot sim-to-real methods Origin and RMA, and ensures better safety during operation. The results indicate that LoopSR enhances locomotion policy capacity and fulfills the purpose of lifelong policy adaptation.
@@ IV-D Ablation Study
We conduct the ablation study to verify each component of LoopSR, including each loss term and the soft update technique. The variants are designed as:
LSR-w/o-con eliminates the contrastive loss in the training process of LoopSR-enc.
LSR-w/o-AE dismisses the reconstruction loss in the training process of LoopSR-enc.
LSR-w/o-su shares the same LoopSR-enc with LSR and excludes the soft update in the adaptation stage.
For LSR-w/o-con and LSR-w/o-AE, we evaluate the prediction accuracy for simulation parameters to test the variants’ ability to reconstruct a more similar simulation. For LSR-w/o-su, we compare the final reward across different terrains to discuss how the strategy in building simulation counterparts matters.
From Tab. IV, LSR with all losses outperforms variants in accurately predicting simulation parameters. The comparable failure of LSR-w/o-con signifies the importance of the contrastive loss [47] in building an interpretable latent space. Besides, LSR also surpasses LSR-w/o-AE with a considerable margin, which proves that the autoencoder structure is beneficial in extracting enough information from the trajectories. Furthermore, as Tab. V shows, LSR-w/o-su drastically underperforms on demanding terrains, demonstrating the necessity of soft update.
Domain | LSR | LSR-w/o-con | LSR-w/o-AE | |
Terrain Accuracy | Stair | |||